Use the normal distributon to the right to answer the questions (a) What percent of the soores are less than 197 b) Out of 1500 randomly selectod scores, about how many would be expected to be greater than 212 (a) The percent of sores tat are less than 19 is[ Round to two decimail places as needed ) b) About soores would be expected to be greater than 21 Round to the nearest whole number as needed ) MacBook Air 8 9 2 O.
Answer #1

a) P(X < 19)

= P((X - \mu )/\sigma < (19 - \mu )/\sigma)

= P(Z < (19 - 21.1)/5.1)

= P(Z < -0.41)

= 0.3409

b) P(X > 21)

= P((X - \mu )/\sigma > (21 - \mu )/\sigma)

= P(Z > (21 - 21.1)/5.1)

= P(Z > -0.02)

= 1 - P(Z < -0.02)

= 1 - 0.4920

= 0.5080

Expected number of scores = 0.5080 * 1500 = 762